# AI Knows Best? The Paradox of Expertise, AI-Reliance, and Performance in Educational Tutoring Decision-Making Tasks

### v1

@@
Abstract
We present an empirical study of how both experienced tutors and non-tutors judge the correctness of tutor praise responses under different Artificial Intelligence (AI)-assisted interfaces types of explanation (textual explanations vs. inline highlighting). We first fine-tuned several Large Language Models (LLMs) to produce binary correctness labels and explanations, achieving up to 88% accuracy and 0.92 F1 score with GPT-4. We then let the GPT-4 models assist 95 participants in tutoring decision-making tasks by offering different types of explanations. Our findings show that although human-AI collaboration outperforms humans alone in evaluating tutor responses, it remains less accurate than AI alone. Moreover, we find that non-tutors tend to follow the AI’s advice more consistently, which boosts their overall accuracy on the task—especially when the AI is correct. In contrast, experienced tutors often override the AI’s correct suggestions and thus miss out on potential gains from the AI’s generally high baseline accuracy. Further analysis reveals that explanations in text reasoning will increase over-reliance and reduce underreliance, while inline highlighting does not. Moreover, neither explanation style actually has a significant effect on performance and costs participants more time to complete the task instead of saving time. Our findings reveal a tension between expertise, explanation design, and efficiency in AI-assisted decision-making, highlighting the need for balanced approaches that foster more effective human-AI collaboration.
1 Introduction
@@
Many educational tasks—such as grading [10, 11], assessment [1], and providing feedback [5, 7, 12]—require educators to make critical judgments about student performance and progress. These tasks are inherently decision-making processes: teachers assess student outputs against learning objectives, deciding how best to score, comment on, or otherwise respond to student work. As educational settings become more complex and class sizes grow, educators often face increasing time and cognitive demands to maintain fair and consistent evaluations [10, 11, 8].
@@
In the present era, Artificial Intelligence (AI) significantly aids human decision-making. Studies show that AI recommendations can improve speed and performance as well as introduce reliance in tasks such as question-answering [18], logical reasoning [2, 16], sentiment analysis [2], and image annotation [13]. Nevertheless, prior work predominantly employed traditional machine-learning methods for non-educational tasks. Little research has examined how the latest Large Language Models (LLMs) could support decision-making in educational contexts, specifically in tutoring. Unlike these tasks, tutoring requires accountability and specialized pedagogical knowledge (e.g., how to give constructive feedback). Given the context-dependent nature of AI’s impact, applying findings from non-educational domains directly to tutoring settings with LLMs poses inherent risks.
@@
These circumstances create a strong incentive to incorporate AI tools that reduce educators’ cognitive load and help them make effective decisions or feedback [12]. For instance, AI-driven feedback could help educators quickly assess how well they are motivating or guiding students. Yet, as discussed above, research on human-AI decision-making in education, especially in tutoring scenarios, remains limited. This paper seeks to close this gap by investigating how LLMs can support decision-making in judging effective tutor responses, focusing on deciding whether certain praise or feedback is appropriate. Specifically, we address the following Research Questions:
@@
RQ1: How accurately can an LLM evaluate the appropriateness of tutor-praising responses, and how does its performance compare to (a) humans working alone and (b) human-AI collaboration?
@@
RQ2: How do differing levels of tutor expertise and various AI explanation styles (e.g., textual reasoning vs. inline highlighting) influence human performance—specifically in terms of accuracy, time efficiency, and reliance on AI?
1.0.1 Our Contribution.
@@
In response to those research questions, we find that an LLM can achieve up to 88% accuracy and 0.92 F1 score in evaluating tutor-praising responses, surpassing both humans alone and human-AI collaboration. We also developed two explanation interfaces and observed that textual reasoning fosters greater reliance on AI; further, humans spend more time when collaborating with AI. Finally, while novices boost their overall accuracy by trusting the AI’s generally correct suggestions (yet over-rely when it is wrong), experienced tutors often override correct AI advice, resulting in under-reliance.
2 Related Work
Prior research on AI-assisted decision-making has produced mixed results, underscoring that its benefits depend on contextual factors. On the one hand, studies indicated that humans are sometimes misled by AI when it makes errors, especially if explanations are overly complex or fail to clarify uncertainties [18, 4, 6]. On the other hand, well-designed explanations can bolster users’ ability to judge the correctness of AI outputs. For instance, [13] found that causal explanations help participants recognize AI unreliability when the AI is wrong, reducing blind reliance. However, if the AI’s decision is correct but the explanation is flawed, most participants do not immediately doubt the overall judgment; rather, they attempt to reconcile the erroneous explanation. Notably, visually misleading cues often carry greater persuasive power than textual errors, and combining incorrect visual and textual explanations can create a “double misdirection” effect [13]. If, however, visual and textual explanations conflict, participants are more likely to question the AI’s decision.
Despite these challenges, well-designed AI systems can complement human decision-makers, achieving performance that neither humans nor AI alone can match [2]. Yet, research also shows that humans frequently either over-rely on AI even when it is wrong or under-rely on it even when AI is correct, often missing the “appropriate reliance” that leverages AI’s suggestion when it is correct and rejecting it when it errs [19, 18, 4, 6]. Even when human-AI teams outperform humans working by themselves, they often fail to exceed AI-only performance. One key to boosting human-AI collaboration performance is enhancing AI explainability. For example, explanation designs that integrate the AI’s reasoning within the question interface (e.g., inline highlights) reduce cognitive load and help users better assess correctness [2, 18]. Additionally, users’ reliance on AI may vary according to their task expertise. For instance, [14] find that non-experts who are less equipped to gauge correctness are more susceptible to over-reliance.
In this study, we aim to explore how different explanation methods, particularly highlighting and text explanations, can help improve human-AI performance in evaluating the effectiveness of tutoring feedback.
2.1 Evaluate Tutor Giving Effective Praise with AI
According to prior work [17], giving effective praise involves genuine and timely acknowledgment of a student’s specific strengths, avoiding overused expressions that lose impact. Rather than saying “good job,” effective praise highlights the student’s accomplishment by connecting it to effort rather than inherent ability. Clear, truthful, and immediate praise fosters greater motivation, resilience, and engagement. Prior work has employed AI to classify tutor responses like using a BERT-based Named Entity Recognition (NER) model to classify the effective praise [12] and obtained up to 73% accuracy and 0.81 in F1 score. They found that the model worked well for effort-related praise but it struggled with outcomes-based praise because of the lack of training data.
We believe that leveraging Large Language Models (LLMs) can effectively address the challenge of insufficient data. As highlighted in the report of [15], LLMs have undergone comprehensive training with vast datasets, enabling them to execute various tasks proficiently. We only need to provide a task description and a few examples [3], and LLMs can effectively judge whether the tutor’s response meets the standards of effective praise. Yet, the random process of text generation [9] makes verifying LLM outputs difficult. As a result, designing explanation methods that clarify the AI’s decision-making is essential for enabling educators to interpret—and, when necessary, override—LLM judgments.
3 Method
3.1 Data Preparation
Through a tutor training system, we collected 216 responses from a simulated scenario where tutors complimented students after solving a math problem. Two senior and experienced tutors and a PhD learning scientist were invited to label the correctness of these responses. The kappa score between the two experienced tutors was 0.76, indicating a substantial consistency, and then the learning scientist resolved the disagreements. This process established our ground truth data. Of the total responses, 155 were labeled as correct (effective praise), and 61 were labeled as incorrect (not effective praise). Among these, four responses were used as a few-shot training data in the prompt, while the rest were used for prediction.
3.2 Prompt Engineering
Table 1: Performance Metrics for Different Models
Model Accuracy Precision Recall F1 Score
gpt-4o-2024-11-20 0.847 0.930 0.852 0.889
gpt-4o-2024-08-06 0.847 0.918 0.865 0.890
gpt-4o-mini-2024-07-18 0.815 0.908 0.826 0.865
gpt-4-1106-preview 0.870 0.885 0.942 0.912
gpt-4-0125-preview 0.861 0.879 0.935 0.906
gpt-4-0613 0.884 0.892 0.955 0.922
gpt-3.5-turbo 0.792 0.917 0.781 0.843
Table 2: Ablation analysis result of the prompt with gpt-4o-2024-08-06
Few-shot Instruction Accuracy Precision Recall F1 Score
X O 0.653 0.976 0.529 0.686
O X 0.824 0.820 0.968 0.888
O O 0.847 0.912 0.871 0.891
We aim to use LLM to assist users in determining if a tutor’s response is effective praise. By providing guidelines for Giving Effective Praise and some selected examples as the prompt, we enable LLM to evaluate whether the tutor’s response meets the criteria for effective praise. Additionally, we ask LLM to generate two explanations for its decision. The first type involves direct text reasoning for the decision. The second type outputs the original input, incorporating HTML tags to Highlight the sections it deems to meet the standards of effective praise. Due to the page limit, the full prompt and example output is at  [^1] https://gist.github.com/anonymousStars/4a56d304515046d806eee4bbfb4d7e2b.
We used this prompt to run through all models in Table 1, the best accuracy of 0.884 is achieved by gpt-4-0613, with an F1 Score of 0.918. Moreover, we further performed a majority vote ensemble on the above models and found that combining the following three models from —”gpt-4o-2024-11-20”, ”gpt-4-0613”, and ”gpt-4-1106-preview”— for majority voting further improved the Accuracy to 0.889 and the F1 Score to 0.924.
We further performed an ablation analysis of the prompt using the gpt-4o-2024-08-06 model. The results, shown in Table 2, indicate that few-shot learning plays a more significant role in this context compared to instructions. However, instructions still contribute by providing additional knowledge that helps enhance the model’s performance and help it create better reasoning and highlight explanations when we observed the output data.
3.3 Interface Design
@@
Figure 1: The five different interfaces we designed for AI-assisted decision-making.
@@
  1. 1.
    N: Task content without AI (No AI control group).
  2. 2.
@@
We use the No AI control group as a baseline to compare users’ performance with and without AI. Moreover, we use the AI label-only control group to compare users’ performance and reliance with different types of AI explanations.
3.4 Participants
@@
3.5 Procedure and Design
Our experiment was approved by the Institutional Review Board (IRB). At the experiment platform, participants started by going through a training module on giving effective praise, which consisted of guidelines and examples explaining why certain tutor’s responses are or are not effective praise. They then proceeded to engage in a task involving five rounds of judging the correctness of Tutor Responses. We employed a counterbalanced design, in which each participant was randomly assigned to one of the possible permutations of the four conditions in Section 3.3, ensuring that each condition occurred in each position equally often across participants. Each condition included 20 tasks.
In addition, the tasks encountered by the participants were also randomized, but we kept these tasks’ distribution and accuracy rates aligned with the original data. That is, in the 20 tasks of each design condition, participants encountered 13 correct tutor responses and 7 incorrect tutor responses. Moreover, to reflect the distribution of real data, we set the AI’s accuracy rate to 0.8, and participants encountered 3 false negatives and 1 false positive at each design condition. Finally, to prevent participants from losing confidence in the AI due to errors, we ensured that participants received correct suggestions from the AI for the first five tasks across all conditions.
3.6 Data Collection and Analyze
We first conducted a 2 x 2 Analysis of Variance (ANOVA) to compare how different groups of participants perform with and without AI. Moreover, we check the participant’s reliance by comparing the reliance rate in No Label condition with other conditions in a one-way ANOVA and post hoc tests. Then, we used repeated measures ANOVA with the presence of Highlighting and Text Reasoning in a 2 x 2 ANOVA to compare participants’ performance and reliance on different types of explanations. Finally, we compared the difference between the experienced Tutor and Prolific participants using an independent sample t-test for the following measures:
3.6.1 Human Performance
3.6.2 Time Spent
We evaluated the average time participants spent on each task by calculating the time interval between their task submissions. When analyzing the time spent, we noted that since the experiment was conducted online, a few participants disconnected midway, engaged in other activities, and then reconnected later, resulting in some extreme completion times. To address this, we applied the Interquartile Range (IQR) method to remove outliers before conducting our analysis, as recommended by [20].
3.6.3 Reliance on AI in Different Conditions
Building on prior research [18, 4, 6, 13], we categorize two forms of reliance on AI as follows according to how participants perform in testing scenarios where the AI provides labels and explanations:
4 Results
4.1 Decision-Making Performance across groups
Figure 2: Correct Percentage across different conditions and groups. Participants from Prolific and working with AI performed better. The error bar is Standard Error.
Table 3: Decision-Making Performance Comparison Between Tutor and Prolific Participants. T-test Results.
@@ 4.1 Decision-Making Performance across groups
Reasoning 0.726 (0.122) 0.748 (0.109) -0.949 0.3450 0.194
Highlighting 0.707 (0.108) 0.762 (0.081) -2.818 0.0059** 0.575
Highlighting + Reasoning 0.729 (0.105) 0.768 (0.076) -2.129 0.0359* 0.435
@@
The results comparing the performance of Tutors and Prolific participants across different task conditions are presented in Figure 2 and Table 3.
We first conducted a 2-way ANOVA to compare how different groups of participants perform with and without AI. The analysis revealed a significant main effect of AI assistance ($F(1,471)=6.83,p=.009,\eta_{p}^{2}=.014$), indicating that participants using AI performed significantly better than those without AI. Moreover, there was a significant main effect of group ($F(1,471)=5.34,p=.021,\eta_{p}^{2}=.011$), indicating that participants from Prolific have better overall performance compared to experienced tutors. The interaction effect between AI assistance and group was not significant ($F(1,471)=0.91,p=.341,\eta_{p}^{2}=.002$). We then conducted a comparison of repeated measures in terms of the presence of Text Reasoning and Inline Highlighting. The results showed no significant main effects or interactions.
When further comparing performance between different groups of users in different conditions, as shown in Table 3, we found that under the conditions of Inline Highlighting, Highlighting + Text Reasoning, and for overall average, the Prolific group significantly outperformed the Tutor group. The Prolific group also marginally significantly outperformed the Tutor group in the AI Label Only case. However, under the conditions without any AI suggestions and with only Text Reasoning explanations, the two groups had no significant difference in performance.
4.2 Time Spent Comparison
Figure 3: Average Time Spent by Condition and Group. The data is obtained after excluding outliers using the 1.5 interquartile range (IQR) method. Participants spend more time on conditions that have AI explanations. The error bar is Standard Error.
Table 4: Comparison of average time spent in seconds after IQR filtering across conditions for Tutor and Prolific participants, including in-significant t-test results.

### v2

@@
Abstract.
We present an empirical study examining how experienced tutors (experts) and non-tutors (novices) evaluate the correctness of tutor praise responses under different AI-assisted decision-support interfaces and explanation styles. We examine human-AI reliance patterns by decomposing interaction errors into over-reliance (accepting incorrect AI suggestions) and under-reliance (rejecting correct AI suggestions), together with time cost as a process-level indicator. Across conditions, human-AI collaboration improved accuracy compared to humans working alone, but consistently underperformed an AI-only baseline, indicating that human judgment introduced additional errors even when assisted by a highly accurate model. Novices benefited more from AI support since they tend to follow AI suggestions, whereas experts frequently overrode correct AI advice, resulting in lower overall performance, revealing a paradox of expertise in educational decision-making. We further compare two explanation modalities: textual reasoning and inline highlighting. Textual reasoning reduced under-reliance when the AI was correct but increased over-reliance when the AI was wrong, while inline highlighting exerted minimal influence on either behavior. Notably, neither explanation modality improved accuracy, and both increased time costs. As a contribution to learning analytics, we demonstrate how reliance patterns (over-reliance and under-reliance) and time cost function as process-level indicators that reveal how users integrate, or fail to integrate, AI recommendations. Our findings underscore the need for adaptive, trust-calibrated explanation strategies in tutor-facing decision support systems that balance accuracy, efficiency, and accountability in human-AI collaboration.
Human-AI Collaboration, Human-AI Interaction, Educational Decision Making, Explainable AI, Tutor-Facing Decision Support
journalyear: 2026copyright: ccconference: LAK26: 16th International Learning Analytics and Knowledge Conference; April 27-May 01, 2026; Bergen, Norwaybooktitle: LAK26: 16th International Learning Analytics and Knowledge Conference (LAK 2026), April 27-May 01, 2026, Bergen, Norwaydoi: 10.1145/3785022.3785035isbn: 979-8-4007-2066-6/2026/04ccs: Human-centered computing Empirical studies in HCIccs: Human-centered computing User interface design
1. Introduction
@@
Many educational tasks, such as grading (armfield2025avalon; chen2025generative), assessment (anderson2003classroom; han2024improving), and providing feedback (thomas2023tutor; hattie2007power; lin2023using; Lin2024; zhao2025slideitright), require educators to make critical judgments about student performance and progress. These tasks are inherently decision-making processes: teachers assess student outputs against learning objectives, deciding how best to score, comment on, or otherwise respond to student work. As educational settings become more complex and class sizes grow, educators often face increasing time and cognitive demands to maintain fair and consistent evaluations (jerrim2020teacher; kreuzfeld2022teachers). When educational assessments are deliberately designed and decomposed into analytic criteria, scoring often reduces to a sequence of binary evidence checks (e.g., present/absent or meets/does not meet) that are subsequently aggregated into partial-credit scores or performance levels. This evidence-centered operationalization is common in learning analytics and teacher-facing decision-support systems, creating clear opportunities for AI to assist in analyzing data and help close the loop from analytics to decision (li2024bringing; cheng2024evidence; clow2012learning; dourado2021teacher; fernandez2024data).
@@
In the present era, AI significantly aids human decision-making. Studies show that AI recommendations can improve speed and performance as well as introduce reliance in tasks such as question-answering (vasconcelos2023explanations), logical reasoning (bansal2021does; si2023large), sentiment analysis (bansal2021does), and image annotation (morrison2023evaluating). Nevertheless, prior work predominantly employed traditional machine-learning methods for non-educational tasks. Little research has examined how the latest Large Language Models (LLMs) could support decision-making in educational contexts, specifically in tutoring. Unlike these tasks, tutoring requires accountability and specialized pedagogical knowledge (e.g., how to give constructive feedback). Given the context-dependent nature of AI’s impact, applying findings from non-educational domains directly to tutoring settings with LLMs poses inherent risks.
@@
These circumstances create a strong incentive to incorporate AI tools that reduce educators’ cognitive load and help them make effective decisions or feedback (lin2023using). For instance, AI-driven feedback could help educators quickly assess how well they are motivating or guiding students. Yet, as discussed above, research on human-AI decision-making in education, especially in tutoring scenarios, remains limited. This paper aims to bridge this gap by examining how LLMs can facilitate effective decision-making in evaluating tutor responses, determining whether certain praise or feedback is suitable. Specifically, we address the following Research Questions:
@@
RQ1: In evaluating the appropriateness of tutor-praise responses, how does performance compare between (a) humans working alone and (b) human–AI collaboration?
@@
RQ2: How do different tutor experience levels and AI explanation styles (textual reasoning vs. inline highlighting) influence human performance, specifically with respect to accuracy, time efficiency, and reliance on AI?
@@
In response to those research questions, we developed two explanation interfaces and observed that textual reasoning fosters greater reliance on AI; further, humans spend more time when collaborating with AI. Finally, while novices boost their overall accuracy by trusting the AI’s generally correct suggestions (yet over-rely when it is wrong), experienced tutors often override correct AI advice, resulting in under-reliance.
2. Related Work
2.1. AI-Assisted Decision Making
Prior research on AI-assisted decision-making shows mixed results: humans may be misled by AI errors, especially with unclear or complex explanations (vasconcelos2023explanations; buccinca2021trust; dietvorst2015algorithm), yet well-designed explanations can help users detect unreliability and avoid blind reliance (morrison2023evaluating). Visual errors often mislead more than textual ones, and conflicting cues may prompt users to question AI (morrison2023evaluating). Overall, humans tend to either over-rely on wrong AI suggestions (accepting errors uncritically) or under-rely on correct AI suggestions (ignoring accurate advice), and thus rarely achieve appropriate reliance, where users trust AI when it is right and reject it when it is wrong (vereschak2021evaluate). Moreover, research show that different explanation designs such as inline highlights can reduce cognitive load (bansal2021does). Reliance also varies by expertise: non-experts are more susceptible to over-reliance (morrison2023impact). In this study, we examine how different explanation styles affect human–AI collaboration in evaluating tutoring feedback for users with different levels of expertise, and discuss their implications for learning analytics in educational contexts.
2.2. Giving Effective Praise in Tutoring and Evaluating it with AI
Prior work in tutor training emphasizes effective praise as feedback that is specific to what the learner did, focusing on process or effort rather than personal traits or outcomes (thomas2023tutor). For example, “You checked each step carefully and revised when something didn’t look right, keep using that approach” praises the learner’s strategy, whereas “You’re so smart!” or “Perfect, you got it right” focuses on inherent ability or just the outcome. This distinction aligns with growth-mindset research, which shows that attributing success to controllable strategies and effort encourages persistence and continued engagement, while ability-focused praise can promote fixed beliefs (blackwell2007implicit). Building on these guidelines, prior work has employed AI to classify tutor responses, such as using a BERT-based Named Entity Recognition (NER) model to classify effective praise (lin2023using), obtaining up to 73% accuracy and a F1 score of 0.81. They found that the model performed well for effort-related praise but struggled with outcomes-based praise due to the lack of training data.
3. Method
@@
Figure 1. The five different interfaces we designed for AI-assisted decision-making.
3.1. Data Preparation
Through a tutor training system, we collected 216 responses from a simulated scenario where tutors complimented students after solving a math problem. Two senior and experienced tutors and a PhD learning scientist were invited to label the correctness of these responses. The kappa score between the two experienced tutors was 0.76, indicating substantial consistency, and then the learning scientist resolved the disagreements. This process established our ground truth data. Of the total responses, 155 were labeled as correct (effective praise), and 61 were labeled as incorrect (not effective praise). Among these, four responses were used as a few-shot training data in the prompt, while the rest were used for prediction.
3.2. Prompt Engineering and Model Selection
We prompted an LLM to assign a binary label and produce the following two explanations: text reasoning and inline highlights (by HTML tags) over the original input. Based on our prior evaluation on this dataset (chen2025identifying), gpt-4-0613 performed best with 0.88 accuracy and 0.92 F1 score, so we use its outputs in the user study. For more about prompt engineering and evaluation on this dataset, please refer to our prior work at (chen2025identifying).
3.3. Interface Design
@@
  1. (1)
    N: Task content without AI (No AI baseline group).
  2. (2)
@@
We use the No AI group as a baseline to compare users’ performance with and without AI. Moreover, we use the AI label-only control group to compare users’ performance and reliance on different types of AI explanations.
3.4. Participants
@@
3.5. Procedure and Design
Our experiment was approved by the Institutional Review Board (IRB). At the experiment platform, participants started by going through a training module on giving effective praise, which consisted of guidelines and examples explaining why certain tutors’ responses are or are not effective praise. They then proceeded to engage in a task involving five rounds of judging the correctness of Tutor Responses. We employed a counterbalanced design, in which each participant was randomly assigned to one of the possible permutations of the five conditions in subsection 3.3, ensuring that each condition occurred in each position equally often across participants. Each condition included 20 tasks.
Moreover, the tasks encountered by the participants were also randomized, but we kept these tasks’ distribution and accuracy rates aligned with the original data. That is, in the 20 tasks of each design condition, participants encountered 13 correct tutor responses and 7 incorrect tutor responses. Moreover, to reflect the distribution of real data, we set the AI’s accuracy rate to 0.8, and participants encountered 3 false negatives and 1 false positive at each design condition. Finally, to prevent participants from losing confidence in the AI due to errors, we ensured that participants received correct suggestions from the AI for the first five tasks across all conditions.
Table 1. Decision-Making Performance Comparison Between Tutor and Prolific Participants. T-test Results.
@@ 3.5. Procedure and Design
Text Reasoning 0.726 (0.122) 0.748 (0.109) -0.949 0.3450 0.194
Inline Highlighting 0.707 (0.108) 0.762 (0.081) -2.818 0.0059** 0.575
Inline Highlighting + Text Reasoning 0.729 (0.105) 0.768 (0.076) -2.129 0.0359* 0.435
@@
Figure 2. Correct Percentage across different conditions and groups. Participants from Prolific and working with AI performed better. The error bar is Standard Error.
3.6. Data Collection and Analysis
We first conducted a 2 x 2 Analysis of Variance (ANOVA) to compare how different groups of participants perform with and without AI label and explanation. Moreover, we used a one-way ANOVA with post hoc tests to compare error rates on AI-correct and AI-incorrect items between the No Label baseline and AI-assisted conditions. Then, we used repeated measures ANOVA with the presence of Highlighting and Text Reasoning in a 2 x 2 ANOVA to compare participants’ performance and reliance on different types of explanations while the label is always presented. Finally, we compared the difference between the experienced Tutor and Prolific participants using an independent sample t-test for the following measures:
3.6.1. Human Performance
3.6.2. Time Spent
We evaluated the average time participants spent on each task by calculating the time interval between their task submissions. When analyzing the time spent, we noted that since the experiment was conducted online, a few participants disconnected midway, engaged in other activities, and then reconnected later, resulting in some extreme completion times. To address this, we applied the Interquartile Range (IQR) method to remove outliers before conducting our analysis, as recommended by (vinutha2018detection).
3.6.3. Reliance on AI in Different Conditions
Building on prior research (vasconcelos2023explanations; buccinca2021trust; dietvorst2015algorithm; morrison2023evaluating), we categorize two forms of reliance on AI as follows, according to how participants perform in testing scenarios where the AI provides labels and explanations:
4. Results
4.1. Decision-Making Performance
The results comparing the performance of Tutors and Prolific participants across different task conditions are presented in Figure 2 and Table 1. We first conducted a 2-way ANOVA to compare how different groups of participants perform with and without AI. The analysis revealed a significant main effect of AI assistance ($F(1,471)=6.83,p=.009,\eta_{p}^{2}=.014$), indicating that participants using AI performed significantly better than those without AI. Moreover, there was a significant main effect of group ($F(1,471)=5.34,p=.021,\eta_{p}^{2}=.011$), indicating that participants from Prolific have better overall performance compared to experienced tutors. The interaction effect between AI assistance and group was not significant ($F(1,471)=0.91,p=.341,\eta_{p}^{2}=.002$).
We then conducted a comparison of repeated measures in terms of the presence of Text Reasoning and Inline Highlighting while the AI label is always present. The results showed no significant main effects or interactions.
Table 2. Comparison of average time spent in seconds after IQR filtering across conditions for Tutor and Prolific participants, including in-significant t-test results.